# Arcee’s MergeKit: A Toolkit for Merging Large Language Models

### v2

@@ 5.1 Practical Example: Applying Model Merging in Medical Domain
As illustrated in Table 1, we experimented with a range of merging techniques available on MergeKit, including Linear intERPolation (LERP), SLERP, TIES, and DARE-TIES, to merge the Meditron-7B [^9] Meditron-7B checkpoint is based on Llama2-7B base model, which is extensively pretrained on a comprehensively curated medical corpus. Chen et al. (2023) checkpoint with Llama2-7B chat model Touvron et al. (2023). Notably, both models are based on the Llama2-7B base model. The evaluation results are depicted in Table 1. According to the findings, all the merged models outperform the Meditron-7B model across various medical benchmarks, including the US Medical License Exam (USMLE) Jin et al. (2021), Medical Multiple-Choice Question Answering (MedMCQA) Pal et al. (2022), and PubMed [^10] https://pubmed.ncbi.nlm.nih.gov/ Question Answering (PubMedQA) Jin et al. (2019). Furthermore, models merged using LERP and SLERP techniques exhibit superior performance over the Llama2-7B chat model in general benchmarks. Our empirical experiments highlight the varying capabilities of merged models and provide comparative performance insights. Within the medical domain, the SLERP method appears to outperform others. However, more importantly, these experiments reveal how model merging can lead to the development of more generalized models with enhanced capabilities across diverse applications.
@@ 6 Conclusion and Future Work
In this paper, we have introduced Mergekit, an innovative open-source tool designed to facilitate the seamless integration of large language models. Beyond detailing the functionalities of the library, we have provided a synthesis of recent literature on model merging from an engineering standpoint, focusing specifically on two primary categories: Merging Models with Identical Architectures and Initializations, and Merging Models with Identical Architectures but Different Initializations. Furthermore, we offered insights on incorporating new merging techniques, thereby encouraging researchers in the open-source community to contribute their novel methods by leveraging our existing capabilities. It is crucial to emphasize that Mergekit represents a dynamic project, committed to the continuous incorporation of new methodologies through collaborative efforts with the open-source community. For an up-to-date list of currently supported methods, we invite readers to consult the ReadMe section of our repository [^11] https://github.com/arcee-ai/mergekit.git.
Ethics Statement

### v3

@@ 4.1 Practical Example: Applying Model Merging in Medical Domain
As illustrated in Table 1, we experimented with a range of merging techniques available in MergeKit, including Linear intERPolation (LERP), SLERP, TIES, and DARE-TIES, to merge Meditron-7B [^9] Meditron-7B checkpoint is based on Llama2-7B base model, which is extensively pretrained on a comprehensively curated medical corpus. Chen et al. (2023) with the Llama2-7B chat model Touvron et al. (2023). Both models are based on the Llama2-7B base model. The evaluation results are depicted in Table 1. According to the findings, all the merged models outperform the Meditron-7B model across various medical benchmarks, including the US Medical License Exam (USMLE) Jin et al. (2021), Medical Multiple-Choice Question Answering (MedMCQA) Pal et al. (2022), and PubMed [^10] https://pubmed.ncbi.nlm.nih.gov/ Question Answering (PubMedQA) Jin et al. (2019). Furthermore, models merged using LERP and SLERP techniques exhibit superior performance over the Llama2-7B chat model in general benchmarks. Our empirical experiments highlight the varying capabilities of merged models and provide comparative performance insights. Within the medical domain, the SLERP method appears to outperform others. However, more importantly, these experiments reveal how model merging can lead to the development of more generalized models with enhanced capabilities across diverse applications.
Recent studies emphasize the importance of merging fine-tuned models into their base models to address challenges like catastrophic forgetting and skill transfer Alexandrov et al. (2024); Siriwardhana et al. (2024). This technique helps maintain prior knowledge while integrating new capabilities. We employed several merging techniques, each with its own hyper-parameters, such as the contribution of each pre-trained model and parameter masking in task vectors.
@@ 5 Conclusion and Future Work
In this paper, we introduce MergeKit, an innovative open-source tool for seamlessly integrating LLMs. We detail its functionalities and provide an overview of recent model merging literature from an engineering perspective. Additionally, we offer insights on incorporating new merging techniques, encouraging community contributions. MergeKit is a dynamic project, committed to continuously integrating new methodologies through collaborative efforts with the open-source community.
Ethical Considerations